The perimeter of a triangle is 24cm. One side is 2cm. longer than the shortest side. The longest side is 2cm less than twice the length of the shorest side. Find the length of each side of the triangle.
Answer by stanbon(75887)

The perimeter of a triangle is 24cm. One side is 2cm. longer than the shortest side. The longest side is 2cm less than twice the length of the shorest side. Find the length of each side of the triangle.
-----------
Let the shortest side be "x".
A 2nd side is "x+2"
The longest side is "2x-2"
EQUATION:
x + x+2 + 2x-2 = 24
4x=24
x=6 cm (shortest side)
x+2 = 8 cm (2nd shde)
2x-2= 10 cm (longest side)
===============
Cheers,
Stan H.